OpenStudy (anonymous):

EXPLAIN IM CONFUSED , PLEASE!! would you make more intrest by investing $5,000 in a 6-month certificate of deposit (CD) at 6% intrest or by investing $3,000 in a 12-month CD at 4% intrest EXPLAIN

OpenStudy (anonymous):

monthly

OpenStudy (cwtan):

Assume it is a geometrical series a=5000 r=1.06 and we want to find the 6th Term 5000(1.06)^6=7092.60 so it gain 2092.60 a=3000 r=1.04 10th Term 3000(1.04^10)=4440.73 so it gain only 1440.73 so the 1st one gain more interest

OpenStudy (cwtan):

Oops it's 12th term for the 2nd one sorry i make a mistake 3000(1.04)^12=4803.10 gain 1803.10